# Rapid formation of a very massive star $>\!\!50000\,\mathrm{M}_{\odot}$ and subsequently an IMBH from runaway collisions

### v1

@@ Abstract
Context. We present simulations of a massive young star cluster using Nbody6++GPU and MOCCA. The cluster is initially more compact than previously published models, with one million stars, a total mass of $5.86\times 10^{5}\leavevmode\nobreak\ \mathrm{M}_{\odot}$, and a half-mass radius of $0.1\leavevmode\nobreak\ \mathrm{pc}$.
@@
Results. We find that runaway stellar collisions in the cluster core produce a VMS exceeding $5\times 10^{4}\leavevmode\nobreak\ \mathrm{M}_{\odot}$ within 5 Myr, which subsequently collapses into an IMBH.
Conclusions. Our model suggests that dense stellar environments may enable the formation of very massive stars and massive black hole seeds through runaway stellar collisions. These results provide a potential pathway for early black hole growth in star clusters and offer theoretical context for interpreting recent JWST observations of young, compact clusters at high redshift.
@@ 1 Introduction
Most galaxies contain central supermassive black holes (SMBH, $10^{6}-10^{9}\leavevmode\nobreak\ \rm M_{\odot}$), embedded in a dense nuclear star cluster. Recent observations revealed the existence of even more massive SMBH ($10^{9}-10^{10}\leavevmode\nobreak\ \rm M_{\odot}$) during the first billion years after the big bang (see Inayoshi et al., 2020 and e.g. Adamo et al., 2024; Harikane et al., 2023; Kokorev et al., 2023, 2024; Goulding et al., 2023; Übler et al., 2023, 2024; Furtak et al., 2024; Matthee et al., 2024; Maiolino et al., 2024; Juodžbalis et al., 2024; Napolitano et al., 2024; Akins et al., 2025). This has sparked renewed interest into the question how such SMBHs could grow in such a relatively short time; if they grow via gas or star accretion from lower mass ”seed” black holes (BHs). Often seed BHs are identified with intermediate mass black holes (IMBH, mass range of $10^{2}-10^{5}\leavevmode\nobreak\ \rm M_{\odot}$, see for a review Greene et al., 2020), but seed BHs may have to be as massive as $10^{6}\leavevmode\nobreak\ \rm M_{\odot}$ to explain the most massive SMBH in case of (super-)Eddington accretion (Shapiro, 2005; Schneider et al., 2023).
1.1 How do seed black holes form?
@@
The formation of central massive BHs from a primordial gas cloud in galactic nuclei has already been postulated by Rees (1984). Formation channels quoted in his famous Fig. 1 are a dense star cluster undergoing stellar collisions and mergers, a dense cluster of BHs, or a single supermassive star. We summarize the current ideas about these channels here, following Inayoshi et al. (2020) and Volonteri et al. (2021):
  1. 1.
@@
    Gravitational mergers between massive stars as a source of IMBH seeds have been discussed in the literature already for decades (Sanders, 1970; Lee, 1987; Quinlan & Shapiro, 1990; Portegies Zwart et al., 1999a, 2004; Giersz et al., 2015; Mapelli, 2016; Reinoso et al., 2021; Wang et al., 2022a; Rantala et al., 2024; Fujii et al., 2024b). This “fast” channel (few Myrs after cluster formation and natal gas expulsion), following Greene et al. (2020), leads to formation of a very massive star (VMS) by mass segregation and subsequent inelastic collisions of stars in the central region of a dense star cluster. We follow the definition of Fricke (1973); Hara (1978); Langbein et al. (1990) that a supermassive star (SMS) is characterized by the non-existence of stable nuclear burning, since the Kelvin-Helmholtz contraction leads directly to relativistic collapse; in the first cited paper $5\cdot 10^{5}\mathrm{M}_{\odot}$ was given as the limit, but the exact number depends on many parameters; our massive object clearly stays below that limit and is therefore denoted as a VMS. One object outgrows the other ones by growing faster through collisions than its stellar evolution lifetime, which is often called ”runaway” growth. Such condition is not necessary for the collisional fast growth of a massive star. Up to now most works use either relatively small particle number for $N$-body simulations or approximate Fokker-Planck or Monte Carlo models. Until recently, computational limitations made $N$-body simulations with large particle numbers ($\gtrsim 10^{6}$) unfeasible. The analysis of observational data of nuclear star clusters (NSCs) and the comparison with relevant timescales, further suggested that the collision time in these systems plays a critical role for the formation of their SMBHs (Escala, 2021). In order to compare the collision timescale with the age of the system, Vergara et al. (2023) derived a critical mass scale for the formation of massive objects.This result is later extended to diverse stellar systems, encompassing various initial conditions, initial mass functions, and evolution scenarios (Vergara et al., 2024). Under the concept of critical mass, Liempi et al. (2025) successfully replicated the observed mass function shape between NSCs and SMBHs with semi-analytic models.
  2. 2.
@@
    Gravitational and general relativistic mergers between stellar-mass BHs can lead to the formation of single massive BHs Zel’dovich & Podurets (1966), an idea followed up by Rees (1984). This scenario has been in recent years re-evaluated and re-discovered under the term “slow” IMBH growth (around 100 Myr to Gyrs). Mass segregation and stellar evolution generates central BH subsystems, but different from the ideas of (Zel’dovich & Podurets, 1966) they are mixed with other stars, and contain binaries, there are not just pure single BH systems. Such models have been shown to generate IMBHs with masses of order $10^{2}\,\mathrm{M}_{\odot}$ to $10^{4}\,\mathrm{M}_{\odot}$(e.g. Portegies Zwart & McMillan, 2002; Miller & Hamilton, 2002; Giersz et al., 2015; Rodriguez et al., 2019; Arca Sedda et al., 2019; Arca-Sedda et al., 2021; Arca Sedda et al., 2023a; Di Carlo et al., 2020a, b; Kremer et al., 2020a; Di Carlo et al., 2021; González et al., 2021; Leveque et al., 2022a; Maliszewski et al., 2022; Askar et al., 2023; Davis et al., 2024; González Prieto et al., 2024). Therefore the evolution towards an IMBH depends sensitively on initial parameters such as central density, binary fraction (e.g. Rizzuto et al., 2021, 2022; Arca-Sedda et al., 2021; Arca Sedda et al., 2021a; Mapelli et al., 2022) and there is not necessarily a phase of relativistic collapse of a BH subsystem (Zel’dovich & Podurets, 1966). The presence of gas inflows, for example after galaxy mergers, can substantially contribute to steepen the gravitational potential and enhance the probability for collisions (Davies et al., 2011). As a result, this channel can potentially contribute substantially to explain the SMBH population (Lupi et al., 2014; Kroupa et al., 2020; Gaete et al., 2024).
  3. 3.
@@
    Direct collapse of extremely massive gas clouds can result in massive BHs of order $10^{4}$ $\mathrm{M}_{\odot}$ to $10^{6}$ $\mathrm{M}_{\odot}$ effectively bypassing all stellar evolution phases (e.g. Bromm & Loeb, 2003; Begelman et al., 2006, 2008a; Begelman, 2010; Latif et al., 2013; Mayer et al., 2015). However, this channel is difficult to establish, as even trace amounts of dust, metals or formation of molecular hydrogen could cause fragmentation (e.g. Omukai et al., 2008; Latif et al., 2014, 2016). The resulting objects are expected to either form supermassive stars or quasi-stars (BHs embedded in gaseous envelopes), ultimately collapsing to form massive BHs (Begelman et al., 2008b; Hosokawa et al., 2013; Schleicher et al., 2013; Haemmerlé, 2020; Kiyuna et al., 2024). We include here also the possibility of stars that form so massive ($10^{5}-10^{6}$ $\mathrm{M}_{\odot}$) that they explode by general-relativistic instability Supernovae (SNe) (e.g. Shibata & Shapiro, 2002; Sakurai et al., 2015; Uchida et al., 2017; Ugolini et al., 2025). Under more realistic direct collapse conditions, one may not expect a simple direct collapse, but has to consider the possibility of fragmentation and subsequent mergers of the clumps (e.g. Boekholt et al., 2018; Tagawa et al., 2020; Schleicher et al., 2022; Reinoso et al., 2023; Rantala et al., 2024; Kritos et al., 2024). The presence of strong far-ultraviolet radiation can allow the formation of stars more massive than $10^{4}$ $\mathrm{M}_{\odot}$ at $Z\lesssim 10^{-3}Z_{\odot}$, while for $Z\simeq 10^{-2}Z_{\odot}$ occurs fragmentation forming dense star cluster that can host stars with $10^{3}$ $\mathrm{M}_{\odot}$ (Chon & Omukai, 2024).
  4. 4.
@@
    Massive Population III (Pop-III) stars have been postulated to produce seed IMBHs with masses of order $10^{2}$ $\mathrm{M}_{\odot}$ through direct collapse above the pair instability mass gap (e.g. Bromm & Larson, 2004; Bromm, 2013; Woosley, 2017; Haemmerlé et al., 2020; Mestichelli et al., 2024). (Extremely massive) Pop-III stars can merge with other Pop-III stars in their host clusters before collapse to produce even more massive IMBHs above the pair-instability mass gap during the fast gravitational runaway merger phase, as outlined above (e.g. Katz et al., 2015; Sakurai et al., 2017; Reinoso et al., 2018, 2021; Tanikawa et al., 2022; Wang et al., 2022a; Mestichelli et al., 2024). Collisions between Pop-III stars, which are also accreting material at rates of $10^{-3}$ $\mathrm{M}_{\odot}\,\mathrm{yr}^{-1}$, can lead to the formation of BHs with masses of the order of $10^{4}$ $\mathrm{M}_{\odot}$ (Reinoso et al., 2025).
@@
Previous work has shown that rapid accretion can lead to very large radii of the supermassive stars forming in such scenarios (up to $100$ AU for $10^{4}$ M stars), particularly when the timescale for mass gain is shorter than the Kelvin-Helmholtz timescale of the stars (Hosokawa et al., 2013; Schleicher et al., 2013; Haemmerlé et al., 2018; Nakauchi et al., 2020). The formation of VMS has been investigated by Nakauchi et al. (2020) with the MESA code (Paxton et al., 2011, 2013, 2015, 2018, 2019) to simulate the evolution of massive stars up to $3000\leavevmode\nobreak\ \rm M_{\odot}$, accounting for stellar wind mass loss from metal-poor to solar metallicity stars. They found that such stars grow from about $30\leavevmode\nobreak\ \rm R_{\odot}$ at the MS to around $10^{4}\leavevmode\nobreak\ \rm R_{\odot}$ in the post-MS phase. Besides Martinet et al. (2023) used the GENEC code (Eggenberger et al., 2008) to model rotating and non-rotating stars of $180$ to $300\leavevmode\nobreak\ \rm M_{\odot}$ with a metallicity of $Z=0.014$, finding MS lifetimes of about $2.2-2.6$ Myr.
1.2 How do the progenitor star clusters form?
@@
Globular cluster formation takes place both in the local universe as well as at high redshifts (Adamo et al., 2024; Mowla et al., 2024, see discussion below). High metallicity clusters may form from high density gas clouds, e.g. in merging galaxies or fragmented tidal arms, and are frequently observed in interacting galaxies. Formation of low metallicity clusters in the early universe, in a cosmological context, is quite difficult regarding the baryonic content of clusters - very narrow metallicity range, which excludes a strong self-enrichment during the formation phase. Recently, Kimm et al. (2016) succeeded, for the first time, in making metal-poor globular clusters at high redshift ($z>7$) during the epoch of reionization in dwarf galaxies in cosmological simulations. These metal-poor clusters will later be accreted into Milky Way - type galaxies. In Kimm et al. (2016), two out of two simulated dwarf galaxies formed a globular cluster, which is more than sufficient to explain the observed population in the Milky Way. It is even conceivable that still denser and/or more massive globular clusters may be formed through this channel considering that they were modeling two randomly chosen halos.
@@
Adamo et al. (2024) found a few compact clusters with intrinsic masses around $10^{6}\rm\leavevmode\nobreak\ \leavevmode\nobreak\ \rm M_{\odot}$ formed $460$ Myr after the Big Bang with the James Webb Space Telescope (JWST). The presence of extremely dense and massive clusters at high redshifts could be essential to explain the high nitrogen-oxygen abundances observed in galaxies (Bouwens et al., 2010; Tacchella et al., 2023; Marques-Chaves et al., 2023; Nagele & Umeda, 2023). VMSs formed in these dense clusters could contribute to these observations through hydrogen-burning nucleosynthesis (Charbonnel et al., 2023). If the VMS is metal-enriched, it can emit powerful winds (Sander & Vink, 2020; Vink, 2022) that can enrich the gas with nitrogen at the expense of oxygen (Glebbeek et al., 2009). Super-compact massive star clusters can form during galaxy mergers (Renaud et al., 2015), which trigger starburst environments and create clusters with tens of thousands to millions of stars. High-redshift, low-metallicity galaxies (Pettini et al., 1997) undergoing mergers are ideal environments for forming massive, compact star clusters where VMS can develop. In conclusion, we are motivated from theory, simulations and observations alike to study the creation of VMSs in dense and massive star clusters that have formed in the early universe.
@@
This paper provides $N$-body models for clusters very similar to the recently observed young massive clusters in the early universe (Mowla et al., 2024; Adamo et al., 2024), and shows how they would form a VMS and an IMBH. It gives also sound support by a full and direct star cluster $N$-body simulation for a long-standing conjecture, that VMSs of more than $10^{4}\mathrm{M}_{\odot}$ do form quickly and early therein by fast (possibly runaway) merging of normal stars with a massive star, as has been predicted by Monte Carlo simulations (cf. Sect. 2.2). Our $N$-body simulations are using Nbody6++GPU, one if not the most advanced and accurate integration code with regard to legacy, astrophysics, numerical and physical accuracy and performance (Aarseth, 1999a; Kamlah et al., 2022a; Spurzem & Kamlah, 2023). This code is able to routinely simulate a globular cluster size star cluster with all necessary astrophysical ingredients (Wang et al., 2015, 2016; Kamlah et al., 2022a; Arca Sedda et al., 2023a). While in this work we leverage our code to simulate an unprecedentedly dense and massive star cluster, the huge computational costs limits the number of realisations to one. The formation and growth of IMBH by mergers with other BHs will be a source of gravitational wave emission (Rizzuto et al., 2021; Arca Sedda et al., 2021b; Rizzuto et al., 2022) and tidal disruption events (Stone & Metzger, 2016; Rizzuto et al., 2023).

### v2

# Rapid formation of a very massive star $50000\,\mathrm{M}_{\odot}$ and subsequently an IMBH from runaway collisions

@@ Abstract
Context. We present simulations of a massive young star cluster using Nbody6++GPU and MOCCA. The cluster is initially more compact than previously published models, with one million stars, a total mass of $5.86\times 10^{5}\penalty 10000\ \mathrm{M}_{\odot}$, and a half-mass radius of $0.1\penalty 10000\ \mathrm{pc}$.
@@
Results. We find that runaway stellar collisions in the cluster core produce a VMS exceeding $5\times 10^{4}\penalty 10000\ \mathrm{M}_{\odot}$ within 5 Myr, which subsequently collapses into an IMBH. We stress that further work on stellar astrophysics is needed, particularly in the context of VMS formation, which currently represents important uncertainties.
Conclusions. Our model suggests that dense stellar environments may enable the formation of very massive stars and massive black hole seeds through runaway stellar collisions. These results provide a potential pathway for early black hole growth in star clusters and offer a theoretical context for interpreting recent JWST observations of young, compact clusters at high redshift.
@@ 1 Introduction
Most galaxies contain central supermassive black holes (SMBH, $10^{6}-10^{9}\penalty 10000\ \rm M_{\odot}$), embedded in a dense nuclear star cluster. Recent observations revealed the existence of even more massive SMBH ($10^{9}-10^{10}\penalty 10000\ \rm M_{\odot}$) during the first billion years after the big bang (see Inayoshi2020 and e.g. Adamo2024; Harikane2023; Kokorev2023; Kokorev2024; Goulding2023; Ubler2023; Ubler2024; Furtak2024; Matthee2024; Maiolino2024; Juodzbalis2024; Napolitano2025; Akins2025). This has sparked renewed interest into the question how such SMBHs could grow in such a relatively short time; if they grow via gas or star accretion from lower mass ”seed” black holes (BHs). Often seed BHs are identified with intermediate mass black holes (IMBH, mass range of $10^{2}-10^{5}\penalty 10000\ \rm M_{\odot}$, see for a review Greene2020), but seed BHs may have to be as massive as $10^{6}\penalty 10000\ \rm M_{\odot}$ to explain the most massive SMBH in case of (super-)Eddington accretion (Shapiro2005; Schneider2023).
@@
The formation of central massive BHs from a primordial gas cloud in galactic nuclei has already been postulated by Rees1984. Formation channels quoted in his famous Fig. 1 are a dense star cluster undergoing stellar collisions and mergers, a dense cluster of BHs, or a single supermassive star. We summarize the current ideas about these channels here, following Inayoshi2020 and Volonteri2021:
@@
Gravitational mergers between massive stars as a source of IMBH seeds have been discussed in the literature already for decades (Sanders1970; Lee1987; Quinlan1990; PortegiesZwart1999; PortegiesZwart2004; Vanbeveren2009; Giersz2015; Mapelli2016; Reinoso2020; Wang2022; Rantala2024; Fujii2024). This “fast” channel (few Myrs after cluster formation and natal gas expulsion), following Greene2020, leads to formation of a very massive star (VMS) by mass segregation and subsequent inelastic collisions of stars in the central region of a dense star cluster. We follow the definition of Fricke1973; Hara1978; Langbein1990 that a supermassive star (SMS) is characterized by the non-existence of stable nuclear burning, since the Kelvin-Helmholtz contraction leads directly to relativistic collapse; in the first cited paper $5\cdot 10^{5}\mathrm{M}_{\odot}$ was given as the limit, but the exact number depends on many parameters; our massive object clearly stays below that limit and is therefore denoted as a VMS. One object outgrows the other ones by growing faster through collisions than its stellar evolution lifetime, which is often called ”runaway” growth. Such condition is not necessary for the collisional fast growth of a massive star. Up to now most works use either relatively small particle number for $N$-body simulations or approximate Fokker-Planck or Monte Carlo models. Until recently, computational limitations made $N$-body simulations with large particle numbers ($\gtrsim 10^{6}$) unfeasible. The analysis of observational data of nuclear star clusters (NSCs) and the comparison with relevant timescales, further suggested that the collision time in these systems plays a critical role for the formation of their SMBHs (Escala2021). In order to compare the collision timescale with the age of the system, Vergara2023 derived a critical mass scale for the formation of massive objects. This result is later extended to diverse stellar systems, encompassing various initial conditions, initial mass functions, and evolution scenarios (Vergara2024). Under the concept of critical mass, Liempi2025 successfully replicated the observed mass function shape between NSCs and SMBHs with semi-analytic models.
@@
Gravitational and general relativistic mergers between stellar-mass BHs can lead to the formation of single massive BHs Zel'dovich1966, an idea followed up by Rees1984. This scenario has been in recent years re-evaluated and re-discovered under the term “slow” IMBH growth (around 100 Myr to Gyrs). Mass segregation and stellar evolution generates central BH subsystems, but different from the ideas of (Zel'dovich1966) they are mixed with other stars, and contain binaries, there are not just pure single BH systems. Such models have been shown to generate IMBHs with masses of order $10^{2}\,\mathrm{M}_{\odot}$ to $10^{4}\,\mathrm{M}_{\odot}$(e.g. PortegiesZwart2002; Miller2002; Giersz2015; Rodriguez2019; ArcaSedda2019; ArcaSedda2021; ArcaSedda2023; DiCarlo2020; DiCarlo2020a; Kremer2020a; DiCarlo2021; Gonza2021; Leveque2022a; Maliszewski2022; Askar2023; Davis2024; Gonza2024). Therefore the evolution towards an IMBH depends sensitively on initial parameters such as central density, binary fraction (e.g. Rizzuto2021; Rizzuto2022; ArcaSedda2021; ArcaSedda2021a; Mapelli2022) and there is not necessarily a phase of relativistic collapse of a BH subsystem (Zel'dovich1966). The presence of gas inflows, for example after galaxy mergers, can substantially contribute to steepen the gravitational potential and enhance the probability for collisions (Davies2011). As a result, this channel can potentially contribute substantially to explain the SMBH population (Lupi2014; Kroupa2020; Gaete2024).
@@
Direct collapse of extremely massive gas clouds can result in massive BHs of order $10^{4}$ $\mathrm{M}_{\odot}$ to $10^{6}$ $\mathrm{M}_{\odot}$ effectively bypassing all stellar evolution phases (e.g. Bromm2003; Begelman2006; Begelman2008; Begelman2010; Latif2013; Mayer2015). However, this channel is difficult to establish, as even trace amounts of dust, ms or formation of molecular hydrogen could cause fragmentation (e.g. Omukai2008; Latif2014; Latif2016). The resulting objects are expected to either form supermassive stars or quasi-stars (BHs embedded in gaseous envelopes), ultimately collapsing to form massive BHs (Begelman2008; Hosokawa2013; Schleicher2013; Haemmerle2020; Kiyuna2024). We include here also the possibility of stars that form so massive ($10^{5}-10^{6}$ $\mathrm{M}_{\odot}$) that they explode by general-relativistic instability Supernovae (SNe) (e.g. Shibata2002; Sakurai2015; Uchida2017; Ugolini2025). Under more realistic direct collapse conditions, one may not expect a simple direct collapse, but has to consider the possibility of fragmentation and subsequent mergers of the clumps (e.g. Boekholt2018; Tagawa2020; Schleicher2022; Reinoso2023; Rantala2024; Kritos2024). The presence of strong far-ultraviolet radiation can allow the formation of stars more massive than $10^{4}$ $\mathrm{M}_{\odot}$ at $Z\lesssim 10^{-3}Z_{\odot}$, while for $Z\simeq 10^{-2}Z_{\odot}$ occurs fragmentation forming dense star cluster that can host stars with $10^{3}$ $\mathrm{M}_{\odot}$ (Chon2025).
@@
Massive Population III (Pop-III) stars have been postulated to produce seed IMBHs with masses of order $10^{2}$ $\mathrm{M}_{\odot}$ through direct collapse above the pair instability mass gap (e.g. Bromm2004; Bromm2013; Woosley2017; Haemmerle2020; Mestichelli2024). (Extremely massive) Pop-III stars can merge with other Pop-III stars in their host clusters before collapse to produce even more massive IMBHs above the pair-instability mass gap during the fast gravitational runaway merger phase, as outlined above (e.g. Katz2015; Sakurai2017; Reinoso2018; Reinoso2020; Tanikawa2022; Wang2022; Mestichelli2024). Collisions between Pop-III stars, which are also accreting material at rates of $10^{-3}$ $\mathrm{M}_{\odot}\,\mathrm{yr}^{-1}$, can lead to the formation of BHs with masses of the order of $10^{4}$ $\mathrm{M}_{\odot}$ (Reinoso2025).
@@
Previous work has shown that rapid accretion can lead to very large radii of the supermassive stars forming in such scenarios (up to $100$ AU for $10^{4}$ M stars), particularly when the timescale for mass gain is shorter than the Kelvin-Helmholtz timescale of the stars (Hosokawa2013; Schleicher2013; Haemmerle2018; Nakauchi2020). The formation of VMS has been investigated by Nakauchi2020 with the MESA code (Paxton2011; Paxton2013; Paxton2015; Paxton2018; Paxton2019) to simulate the evolution of massive stars up to $3000\penalty 10000\ \rm M_{\odot}$, accounting for stellar wind mass loss from m-poor to solar metallicity stars. They found that such stars grow from about $30\penalty 10000\ \rm R_{\odot}$ at the MS to around $10^{4}\penalty 10000\ \rm R_{\odot}$ in the post-MS phase. Besides Martinet2023 used the GENEC code (Eggenberger2008) to model rotating and non-rotating stars of $180$ to $300\penalty 10000\ \rm M_{\odot}$ with a metallicity of $Z=0.014$, finding MS lifetimes of about 2.22.62.2-2.6 Myr.
@@
Globular cluster formation takes place both in the local universe as well as at high redshifts (Adamo2024; Mowla2024, see discussion below). High metallicity clusters may form from high density gas clouds, e.g. in merging galaxies or fragmented tidal arms, and are frequently observed in interacting galaxies. Formation of low metallicity clusters in the early universe, in a cosmological context, is quite difficult regarding the baryonic content of clusters - very narrow metallicity range, which excludes a strong self-enrichment during the formation phase. Recently, Kimm2016 succeeded, for the first time, in making m-poor globular clusters at high redshift ($z>7$) during the epoch of reionization in dwarf galaxies in cosmological simulations. These m-poor clusters will later be accreted into Milky Way - type galaxies. In Kimm2016, two out of two simulated dwarf galaxies formed a globular cluster, which is more than sufficient to explain the observed population in the Milky Way. It is even conceivable that still denser and/or more massive globular clusters may be formed through this channel considering that they were modeling two randomly chosen halos.
@@
Adamo2024 found a few compact clusters with intrinsic masses around $10^{6}\rm\penalty 10000\ \penalty 10000\ \rm M_{\odot}$ formed $460$ Myr after the Big Bang with the James Webb Space Telescope (JWST). The presence of extremely dense and massive clusters at high redshifts could be essential to explain the high nitrogen-oxygen abundances observed in galaxies (Bouwens2010; Tacchella2023; Marques-Chaves2024; Nagele2023). VMSs formed in these dense clusters could contribute to these observations through hydrogen-burning nucleosynthesis (Charbonnel2023). If the VMS is m-enriched, it can emit powerful winds (Sander2020; Vink2022) that can enrich the gas with nitrogen at the expense of oxygen (Glebbeek2009). Super-compact massive star clusters can form during galaxy mergers (Renaud2015), which trigger starburst environments and create clusters with tens of thousands to millions of stars. High-redshift, low-metallicity galaxies (Pettini1997) undergoing mergers are ideal environments for forming massive, compact star clusters where VMS can develop. In conclusion, we are motivated from theory, simulations and observations alike to study the creation of VMSs in dense and massive star clusters that have formed in the early universe. We note that in the local Universe, the most massive stars detected have masses of approximately $100-300\penalty 10000\ \rm M_{\odot}$ (Doran2013; Keszthelyi2025). Beyond the local Universe, stars are too faint to measure their masses directly with current technology.
@@
This paper provides $N$-body models for clusters very similar to the recently observed young massive clusters in the early universe (Mowla2024; Adamo2024), and shows how they would form a VMS and an IMBH. It gives also sound support by a full and direct star cluster $N$-body simulation for a long-standing conjecture, that VMSs of more than $10^{4}\mathrm{M}_{\odot}$ do form quickly and early therein by fast (possibly runaway) merging of normal stars with a massive star, as has been predicted by Monte Carlo simulations (cf. Sect. 2.2). Our $N$-body simulations are using Nbody6++GPU, one if not the most advanced and accurate integration code with regard to legacy, astrophysics, numerical and physical accuracy and performance (Aarseth1999; Kamlah2022a; Spurzem2023). This code is able to routinely simulate a globular cluster size star cluster with all necessary astrophysical ingredients (Wang2015; Wang2016; Kamlah2022a; ArcaSedda2023). While in this work we leverage our code to simulate an unprecedentedly dense and massive star cluster, the huge computational costs limits the number of realisations to one. The formation and growth of IMBH by mergers with other BHs will be a source of gravitational wave emission (Rizzuto2021; ArcaSedda2021c; Rizzuto2022) and tidal disruption events (Stone2016; Rizzuto2023).